# Quantifying Statistical Significance in Diffusion-Based Anomaly Localization via Selective Inference

### v4

@@ Appendix B Calculating the subinterval $\mathcal{Z}^{\text{sub}}$ for diffusion models
We now show that a reconstruction error $\mathcal{E}$ via diffusion models can be expressed as a piecewise-linear function of $\bm{X}$. To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below. It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$. To make the reverse process a piecewise-linear function, we employ a U-Net composed of piecewise-linear components such as ReLU activation functions and pooling layers. Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$. Moreover, since $f^{(t)}_{\theta}(\mathbf{x}_{t})$ in (4) is a composite function of $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$, it is also a piecewise-linear function. By combining them together, we see that $\mathbf{x}_{t-1}$ is written as a piecewise-linear function of $\mathbf{x}_{t}$. Therefore, the entire reconstruction process is a piecewise-linear function since it just repeats the above operation multiple times (see Algorithm 1). As a result, the entire mapping ${\mathcal{D}}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$. Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=\mathcal{E}(\bm{X}))$ is piecewise-linear. Exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed. The following theorem tells that the subinterval $\mathcal{Z}^{\text{sub}}(\bm{a}+\bm{b}z)$ can be computed by solving a set of linear inequalities.
@@
where $\kappa^{(k^{\prime})}_{i}\in\mathbb{R}$ and $\rho^{(k^{\prime})}_{i}\in\mathbb{R}$ for $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ are the coefficient and the constant of the $k^{\prime}$-th interval $[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, and $K^{{}^{\prime}}_{i}$ is the number of linear pieces of $\mathcal{A}_{i}$. For each $i\in[n]$, there exists $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ such that $z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
@@ Appendix C Comparison methods for numerical experiments
  • proposed: The proposed method uses the parametric programming.
  • w/o-pp: The proposed method use over-conditioning (without parametric programming). The $p$-value is calculated as
@@
  • bonferroni: To control the type I error rate, this method applies the Bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
@@
This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
@@ Appendix E Accelerated reverse processes
Methods for accelerating the reverse process have been proposed in DDPM, DDIM Song et al. (2022). When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$. In this case, equations (3) and (5) can be rewritten as
@@
Therefore, piecewise-linearity is preserved, making the proposed method DMAD-test applicable.
Appendix F Robustness of the proposed method
To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$. We considered the following non-Gaussian distributions with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
  • Skew normal distribution family (SND).
  • Exponentially modified gaussian distribution family (EMG).
  • Generalized normal distribution family (GND) with a shape parameter $\beta$. This distribution family can be steeper than the normal distribution (i.e., $\beta<2$).
  • Student’s $t$-distribution family ($t$-distribution).
Note that these distributions are standardized in the experiments. Figure 7 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$. The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$. Figure 8 shows the results of the robustness experiments.
DMAD-test maintains good performance on the type I error rate for all the considered distribution families.
Figure 7: Non-Gaussian distributions with $d=0.04$
(a) SND
(b) EMG
(c) GND
(d) $t$-distribution
Figure 8: Type I Error Rate for Non-Gaussian distribution families
Appendix G Experimental settings for the MVTec AD dataset
G.1 Experimental settings of Brain Trumor Segmentation 2023 dataset
@@
We evaluate our method on T2-FLAIR MRI brain scans from the Brain Trumor Segmentation 2023 dataset Karargyris et al. (2023); LaBella et al. (2023). T2-FLAIR MRI which comprises 934 non-skull-stripped 3D scans with dimensions of $240\!\times\!240\!\times\!155$. From these scans, we extracted 2D $240\!\times\!240$ axial slices at axis $95$, resized them to $64\!\times\!64$ pixels, and categorized them based on the anomaly annotations into $532$ normal images (without tumors) and $402$ abnormal images (with tumors). For each scan, we estimated the mean and variance from pixel values excluding both the non-brain regions and tumor regions identified in the ground truth, followed by standardization. We randomly selected $312$ normal images for model training. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.6$ and the kernel size of the averaging filter to $3$. Note that, when testing images of the MRI brain scans, the non-brain regions are not treated as anomalous regions $\mathcal{M}_{\bm{X}}$.
G.2 Experimental settings of Brain Trumor Segmentation 2023 dataset
@@
We evaluate our method on the MVTec AD dataset Bergmann et al. (2019), which consists of $15$ object categories. Each category provides a training set of normal images and a test set containing both normal and abnormal images, with image resolutions ranging from $900\times 900$ to $1024\times 1024$ pixels. For our experiments, we select six categories (bottle, cable, grid) and resize all images to $128\times 128$ pixels. For the type I error rate experiments, we randomly select $50$ normal images from each category, and for the power experiments, we select $50$ abnormal images per category. The diffusion model is trained on the remaining normal images with $T=1000$ total diffusion steps, of which the first $T^{\prime}=300$ steps are used for reconstruction using $4$ sampling steps. We apply an averaging filter with a kernel size of $3$. We set the anomaly threshold $\lambda$ to $1.0$ for bottle and to $1.2$ for cable and grid. To demonstrate in the power experiments, we compute the intersection between the anomalous region detected by the diffusion model and the anomaly annotation. Since the images in the MVTec AD dataset are RGB, we redefine the image data as $\tilde{\bm{X}}\in\mathbb{R}^{hw\times 3}$, where $h$ and $w$ denote the image height and width. We then vectorize each image by
$\bm{X}$ $=\mathrm{vec}(\tilde{\bm{X}})$ (38)
$=(X_{1,1},X_{1,2},X_{1,3},\,X_{2,1},X_{2,2},X_{2,3},\dots,X_{hw,1},X_{hw,2},X_{hw,3})^{\top}\in\mathbb{R}^{n},$
@@
$\bm{X}^{\mathrm{ref}}$ $=\mathrm{vec}(\tilde{\bm{X}}^{\mathrm{ref}})$ (39)
$=(\tilde{X}_{1,1},\tilde{X}_{1,2},\tilde{X}_{1,3},\,\tilde{X}_{2,1},\tilde{X}_{2,2},\tilde{X}_{2,3},\dots,\tilde{X}^{\mathrm{ref}}_{hw,1},\tilde{X}^{\mathrm{ref}}_{hw,2},\tilde{X}^{\mathrm{ref}}_{hw,3})^{\top}\in\mathbb{R}^{n}$
@@
With this setup, the our method can be applied in the same way as in §3.
Appendix H Experimental results for the real-world datasets
In this section, we show the experimental results for the real-world datasets. We applied the proposed proposed to the MRI images and MVTec AD dataset and compared with the naive method.
$p_{\rm naive}:0.031$, $p_{\rm selective}:0.266$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.331$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.766$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.360$ Refer to caption $p_{\rm naive}:0.031$, $p_{\rm selective}:0.646$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.753$ Refer to caption $p_{\rm naive}:0.011$, $p_{\rm selective}:0.768$
MRI images without tumors
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.026$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.012$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.007$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.003$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.002$ Refer to caption $p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
MRI images with tumors
Figure 9: An example of the results for applying the proposed proposed and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to MRI images.
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.907$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.010$
(a) Bottle (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.567$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.000$
(b) Cable (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.028$, $p_{\rm selective}:0.789$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
(c) Grid (Left: Normal, Right: Anomaly)
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.769$
$p_{\rm naive}:0.000$, $p_{\rm selective}:0.018$
(d) Transistor (Left: Normal, Right: Anomaly)
Figure 10: An example of applying the proposed method and the naive method to six categories of the MVTec AD dataset (Bottle, Cable, Grid, Transistor). For each category, the left figure shows a normal image and the right figure shows an anomalous image. The proposed $p_{\rm selective}$ remains high for normal samples (True Negatives) and low for anomalous samples (True Positives), demonstrating accurate control of the false detection rate. In contrast, the $p_{\rm naive}$ yields low $p$-values across both normal and anomalous images, indicating an inflated false positive rate and invalidity of the naive method.

### v5

@@ Appendix B Calculating the subinterval $\mathcal{Z}^{\text{sub}}$ for diffusion models
We show that a reconstruction error $\mathcal{E}$ via diffusion models can be expressed as a piecewise-linear function of $\bm{X}$. To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below. It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$. To make the reverse process a piecewise-linear function, we employ a U-Net composed of piecewise-linear components such as ReLU activation functions and pooling layers. Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$. Moreover, since $f^{(t)}_{\theta}(\mathbf{x}_{t})$ in (4) is a composite function of $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$, it is also a piecewise-linear function. By combining them together, we see that $\mathbf{x}_{t-1}$ is written as a piecewise-linear function of $\mathbf{x}_{t}$. Therefore, the entire reconstruction process is a piecewise-linear function since it just repeats the above operation multiple times (see Algorithm 1). As a result, the entire mapping $\mathcal{D}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$. Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=\mathcal{E}(\bm{X}))$ is piecewise-linear. Exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\mathrm{sub}}$ can be computed. The following theorem states that the subinterval $\mathcal{Z}^{\mathrm{sub}}(\bm{a}+\bm{b}z)$ can be computed by solving a set of linear inequalities.
@@
where $\kappa^{(k^{\prime})}_{i}\in\mathbb{R}$ and $\rho^{(k^{\prime})}_{i}\in\mathbb{R}$ for $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ are the coefficient and the constant of the $k^{\prime}$-th interval $L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}$], and $K^{{}^{\prime}}_{i}$ is the number of linear pieces of $\mathcal{A}_{i}$. For each $i\in[n]$, there exists $k^{\prime}\in[K^{{}^{\prime}}_{i}]$ such that $z\in[L^{(k^{\prime})}_{i},U^{(k^{\prime})}_{i}$], then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
@@ Appendix C Comparison methods for numerical experiments
  • proposed: The proposed method uses parametric programming.
  • w/o-pp: The proposed method uses over-conditioning (without parametric programming). The $p$-value is calculated as
@@
  • bonferroni: To control the type I error rate, this method applies the Bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as
@@ Appendix E Accelerated reverse processes
Methods for accelerating the reverse process have been proposed in DDPM, DDIM (Song et al., 2022). When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$. In this case, equations (3) and (5) can be rewritten as
@@
Therefore, piecewise-linearity is preserved, making the proposed method DAL-Test applicable.
Appendix F Experimental settings for the real-world datasets
F.1 Experimental settings of Brain Tumor Segmentation 2023 dataset
@@
We evaluate our method on T2-FLAIR MRI brain scans from the Brain Tumor Segmentation 2023 dataset (Karargyris et al., 2023; LaBella et al., 2023). T2-FLAIR MRI comprises 934 non-skull-stripped 3D scans with dimensions of $240\!\times\!240\!\times\!155$. From these scans, we extracted 2D $240\!\times\!240$ axial slices at axis $95$, resized them to $64\!\times\!64$ pixels, and categorized them based on the anomaly annotations into $532$ normal images (without tumors) and $402$ abnormal images (with tumors). For each scan, we estimated the mean and variance from pixel values excluding both the non-brain regions and tumor regions identified in the ground truth, followed by standardization. We randomly selected $312$ normal images for model training. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.6$ and the kernel size of the averaging filter to $3$. Note that, when testing images of the MRI brain scans, the non-brain regions are not treated as anomalous regions $\mathcal{M}_{\bm{X}}$.
F.2 Experimental settings of MVTec AD dataset
@@
We evaluate our method on the MVTec AD dataset (Bergmann et al., 2019), which consists of $15$ object categories. Each category provides a training set of normal images and a test set containing both normal and abnormal images, with image resolutions ranging from $900\times 900$ to $1024\times 1024$ pixels. For our experiments, we select four categories (bottle, cable, grid, transistor) and resize all images to $128\times 128$ pixels. For the type I error rate experiments, we randomly select $50$ normal images from each category, and for the power experiments, we select $50$ abnormal images per category. The diffusion model is trained on the remaining normal images with $T=1000$ total diffusion steps, of which the first $T^{\prime}=300$ steps are used for reconstruction using $4$ sampling steps. We apply an averaging filter with a kernel size of $3$. We set the anomaly threshold $\lambda$ to $1.0$ for bottle and to $1.2$ for cable and grid. In the power experiments, we compute the intersection between the anomalous region detected by the diffusion model and the anomaly annotation. Since the images in the MVTec AD dataset are RGB, we redefine the image data as $\tilde{\bm{X}}\in\mathbb{R}^{hw\times 3}$, where $h$ and $w$ denote the image height and width. We then vectorize each image by
$\bm{X}=\mathrm{vec}(\tilde{\bm{X}})=(X_{1,1},X_{1,2},X_{1,3},\,X_{2,1},X_{2,2},X_{2,3},\dots,X_{hw,1},X_{hw,2},X_{hw,3})^{\top}\in\mathbb{R}^{n},$ (38)
@@
$\bm{X}^{\mathrm{ref}}=\mathrm{vec}(\tilde{\bm{X}}^{\mathrm{ref}})=(\tilde{X}_{1,1},\tilde{X}_{1,2},\tilde{X}_{1,3},\,\tilde{X}_{2,1},\tilde{X}_{2,2},\tilde{X}_{2,3},\dots,\tilde{X}^{\mathrm{ref}}_{hw,1},\tilde{X}^{\mathrm{ref}}_{hw,2},\tilde{X}^{\mathrm{ref}}_{hw,3})^{\top}\in\mathbb{R}^{n}$ (39)
@@
With this setup, our method can be applied in the same way as in §3.